In tests of a computer component, it is found that the mean time between failures is 520 hours. A modification is made which is supposed to increase the time between failures. Tests on a random sample of 10 modified components resulted in the following times (in hours) between failures. 518 548 561 523 536 499 538 557 528 563 At the 0.05 significance level, test the claim that for the modified components, the mean time between failures is not equal to 520 hours.

=
Answer #1

Solution:

x x2
518 268324
548 300304
561 314721
523 273529
536 287296
499 249001
538 289444
557 310249
528 278784
563 316969
∑x=5371 ∑x2=2888621


Mean ˉx=∑xn

=518+548+561+523+536+499+538+557+528+563/10

=5371/10

=537.1

Sample Standard deviation S=√∑x2-(∑x)2nn-1

=√2888621-(5371)210/9

=√2888621-2884764.19

=√3856.9/9

=√428.5444

=20.7013

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu  = 520

Ha : \mu    \neq 520

Test statistic = t

= (T - \mu ) / S / \sqrt n

= (537-520) / 20.70 / \sqrt 10

= 2.597

Test statistic = t = 2.597

P-value =0.0289

\alpha = 0.05  

P-value < \alpha

0.0289 < 0.05

Reject the null hypothesis .

There is sufficient evidence to suggest that